Solve by the substitution method.Type an ordered pair,the solution of the system is???
4x + 3y = -14
x = 5 - 5y
Help please.

Answer by oscargut(2103)

4x + 3y = -14
x = 5 - 5y
so if you substitute x at the 1st eq then
4(5-5y)+3y=-14
-20y+20+3y=-14
-17y=-34
y=2
then
x=5-5(2)=-5
Answer:solution is (-5,2)